If O, G, H are the circumcentre, centroid , orthocentre of a triangle ABC, then OA + OB + OC =?

Grade:11

1 Answers

Arun
5 years ago
we know that HG =2GO  where G is the centroid of trianglelet a point D between B and CSo, OD =(OB+OC )/2now OA + OB + OC = OA + 2ODwe know that G devide the point A and mib point of apposite side(D) in ratiio  2:1So, OG =(OA 2OD)/(2+1)So, OA + OB + OC =3OG=2OG + OG=GH'+OG =OH